Disney Company

Beta 0.95

Expected return of the market 10.36%

Risk free rate 3.14%

Tax rate 12.02%

Pre-tax cost of debt 3.07%

Equity 157215

Preferred equity 0

Short-term debt 5992

Long-term debt 17681

Total Capital 180888

Fed Rate Increase.

An increase in Federal Reserve interest rate by 200

basis points (2.0%) lifts the borrowing costs of every debt issuer. What

is Disney’s adjusted WACC?
Answer #1

Beta = 0.95, Market Return = 10.36 % and Risk-Free Rate = 3.14 %

Cost of Equity = ke = 3.14 + 0.95 x (10.36 - 3.14) = 9.999 %

Pre-Tax Cost of Debt = 3.07 % and Increase in Issuer Rate owing to Fed Rate increase = 2 %

Adjusted Pre-Tax Cost of Debt = 3.07 + 2 = 5.07 %

Equity = 157215, Short-Term Debt = 5992 and Long-Term Debt = 17681

Total Debt = 5992 + 17681 = 23673

Equity Proportion = [157215 / (157215 + 23673)] = 0.86913 and Debt Proportion = 1-0.86913 = 0.13087

Tax Rate = 12.02 %

Adjusted WACC = (1-Tax Rate) x 5.07 x 0.13087 + 9.999 x 0.86913 = 9.27418 % ~ 9.27 %